OpenStudy (anonymous):

Solve the following equation to get y by itself: 7x-y=20

OpenStudy (anonymous):

What is the first step to getting y by itself?

OpenStudy (anonymous):

Do you want the problem to be worked out step by step?

OpenStudy (anonymous):

no i just need the answer

jigglypuff314 (jigglypuff314):

to get y by itself for 7x-y=20 first add y to both sides to make it positive +y +y so 7x = 20 + y then subtract 20 from both sides -20 -20 7x - 20 = y